OpenStudy (anonymous):

y varies directly with x. When y = 20, x = 50. Find x when y = 36.

OpenStudy (blurbendy):

y = kx 20 = k50 k = 20/50 = 2/5 36 = 2/5*x solve for x

OpenStudy (anonymous):

i still dont understand how to solve from this point